# A Comprehensive Survey of Machine Unlearning Techniques for Large Language Models

### v1

@@ 3.3 Leveraging Auxiliary Models
$\small\min_{\theta_{a}}\mathcal{L}(\theta_{a})=\min_{\theta_{a}}\{\mathcal{L}_{f}(\theta_{a})-\beta\mathcal{L}_{r}(\theta_{a})\},$ (11)
@@
$\small l(y|x;\theta_{ori})-\alpha\cdot{ReLU}(l(y|x;\theta_{a})-l(y|x;\theta_{ori}))$ (12)
@@
Task Vector
Task vectors, defined as $\tau=\theta_{a}-\theta_{ori}$, steer the model behavior by editing the weight space, thus enabling operations such as negation and addition for applications such as unlearning and multi-task learning. Negating task vectors effectively suppress behaviors such as mitigating toxic language generation Ilharco et al. (2023). Liu et al. [2024c] enhanced fine-tuning with modules targeting harmful knowledge: guided distortion, random disassociation, and preservation divergence. Ethos Gao et al. (2024) distinguishes general and undesired knowledge by projecting task vectors onto principal components. By negating only undesired components, Ethos minimizes collateral damage to model utility, thus achieving unlearning.
@@ 3.4 Input/Output-Based Unlearning
Brief Summary
Table 3.2 summarizes the existing unlearning methodologies, their features, representative work, and applications, highlighting challenges in forget quality, utility preservation, and efficiency based on their principles and empirical findings. Methods such as gradient ascent, preference optimization, and reinforcement learning often struggle with utility preservation due to optimization conflicts (e.g., maximizing cross-entropy loss), requiring regularization to mitigate the performance loss. Relabeling-based fine-tuning and knowledge distillation have been reported to suffer from knowledge relearning. This issue arises when models are fine-tuned using small amounts of related or even unrelated samples, or through in-context learning Lynch et al. (2024). Representation engineering methods have been reported to be more susceptible to adversarial attacks, failing to robustly erase knowledge Łucki et al. (2024). Contrastive decoding and input/output-based unlearning are controversial because they do not truly remove knowledge from the models. Task vectors, limited by imprecision in localizing knowledge, are mainly used in AI alignment. From an efficiency standpoint, reinforcement learning and locate-then-unlearn approaches involve higher algorithmic complexity. Reinforcement learning requires training high-quality models to generate reward signals, while locate-then-unlearn methods rely on precise knowledge localization. Moreover, contrastive decoding and input/output-based methods increase the computational overhead during inference, which leads to slower generation.
@@ 4 Benchmarks
This section provides a detailed description of the commonly used benchmarks, addressing areas such as copyright, privacy, and AI alignment. Notably, WMDP and RWKU are designed for the direct-unlearning, while other benchmarks are used within the fine-tuning-then-unlearning paradigm.
@@ 4.1 Unimodal Benchmarks
Who is Harry Potter (WHP) Eldan et al. (2024) evaluates unlearning of Harry Potter-related information using a dataset combining the original books (2.1M tokens) with synthetic content (1M tokens). Unlearning effectiveness is assessed through 300 Harry Potter-related questions scored with a GPT-4-based evaluation. It explores the potential of unlearning in copyright protection.
TOFU Maini et al. (2024) includes 200 synthetic author profiles with 20 question-answer examples each, ensuring no overlap with existing training data. In addition to offering three forget-retain splits (90-10, 95-5, and 99-1), the benchmark includes 100 real-world author profiles and 117 world facts, providing a comprehensive evaluation of model utility after unlearning.
WMDP Li et al. (2024a) comprises 3,668 multiple-choice questions on biosecurity, cybersecurity, and chemical security, curated by experts to evaluate hazardous knowledge while excluding sensitive information. It serves as a benchmark for assessing LLMs’ hazardous knowledge and unlearning methods for AI alignment.
RWKU Jin et al. (2024) includes 200 unlearning targets focused on well-known public figures, with a total of 13,131 multi-level forget probes. These probes are divided into 3,268 fill-in-the-blank probes, 2,879 question-answer probes, and 6,984 adversarial attack probes.
@@ 5 Evaluation Measures
This section outlines common evaluation measures used to access both utility preservation and forgetting quality.
5.1 Utility Preservation Evaluation
We categorize the methods for evaluating model utility into general utility and localized utility. General utility evaluates the model’s overall knowledge, including downstream task performance and related metrics. In contrast, localized utility examines fine-grained changes, such as output or logit probability similarity, in the model’s output before and after unlearning on task-specific data.
5.1.1 General Utility
Accuracy on General Benchmarks
Unlearned models must retain their original world knowledge without degradation during the unlearning process. It is common to use additional benchmarks Joshi et al. (2017); Fu et al. (2023) to evaluate the utility preservation.
GPT-Eval
LLM-as-a-Judge has been widely adopted across various evaluation scenarios due to its flexibility and ability to assess multiple dimensions. It provides semantic insights that traditional measures might overlook. For instance, Ma et al. [2024] leverage GPT-4o Mini to evaluate unlearning algorithms based on correctness, helpfulness, and relevance, assigning an overall score on a scale from 0 to 1.
5.1.2 Localized Utility
ROUGE
is a measure for evaluating word-level alignment between a model’s output and the ground truth answer. The ROUGE-L recall score is used to quantify the overlap between the generated output and the reference answer Maini et al. (2024); Yuan et al. (2024).
Probability
Maini et al. (2024) evaluates the model’s ability to predict the correct answer for a given question by computing the normalized conditional probability of the answer, defined as $P(y|x)=\frac{1}{T}\sum_{t=1}^{T}p(y_{t}|y_{<t}\circ x;\theta_{unl})$.
Truth Ratio
evaluates the relative likelihood of a correct answer compared to an incorrect one for a given question Maini et al. (2024). It is defined as the ratio of the average normalized conditional probability of perturbed incorrect answers $\bar{y}$ from the $\mathcal{A}$, to that of paraphrased correct answer $\tilde{y}$ of the ground truth $y$. Ideally, when a model lacks the relevant knowledge, it should assign similar prediction probabilities to both correct and incorrect answers. Formally, the truth ratio on the forget set is defined as $R_{truth}(y)=\frac{\frac{1}{\mathcal{A}}\sum_{\bar{y}\in\mathcal{A}}p(\bar{y}|x;\theta_{unl})^{1/|\bar{y}|}}{p(\tilde{y}|x;\theta_{unl})^{1/|\tilde{y}|}}$, where $\bar{y}$ and $\tilde{y}$ can be generated by LLMs. It is expressed as $\max(0,1-R_{truth}(y))$ on the retain set and world knowledge datasets.
5.2 Forget Efficacy Evaluation
Utility-based evaluation, which compares the utility changes of a model on the forget set before and after unlearning, is the most commonly assessment method. However, recent studies Lynch et al. (2024); Ma et al. (2024) have highlighted the critical need to differentiate between real forgetting and mere suppression, where the suppressed knowledge can often be recovered through adversarial techniques. This section reviews existing methods for comprehensive forgetting quality evaluation beyond utility-based evaluation.
5.2.1 Input-Based Evaluation
Paraphrasing and other Languages
Using generative models such as GPT, paraphrased questions can be easily generated to test the unlearning robustness Maini et al. (2024). Additionally, inspired by the performance disparity of large models across different languages, Lynch et al. [2024] evaluates the robustness of LLM unlearning by translating inputs into Spanish and Russian.
In-Context Relearning
This method leverages demonstrations to interfere with the model’s generation process, prompting it to answer questions it is supposed to forget rather than refusing them. Lynch et al. [2024] show that providing unlearned models with small amounts of general context related to Harry Potter can effectively resurface suppressed knowledge.
Jailbreak Prompts
Jailbreak methods manipulate LLM behavior using black-box techniques like role-playing and obfuscation or white-box strategies that optimize prefixes and suffixes to maximize affirmative response likelihood. In the context of learning, Lynch et al. [2024] evaluated two jailbreak prompts that successfully access the model’s memory of the target dataset through privilege escalation and role-playing.  Łucki et al. [2024] introduced the improved GCG to bypass the unlearning mechanism by optimizing prompt prefixes and leveraging early token loss.
5.2.2 Logit-Based Evaluation
Kolmogorov-Smirnov Test (KS-Test) is a non-parametric statistical method widely used to compare two probability distributions. In the context of forget quality evaluation, the KS-Test assesses the divergence between the distributions of model outputs (e.g., token probabilities or embeddings) before and after unlearning specific data. By computing the maximum absolute difference between the cumulative distribution functions (CDFs) of the two distributions, the test quantifies the extent to which the influence of the unlearned data has been mitigated. A smaller KS statistic indicates a higher degree of successful unlearning. The test’s non-parametric nature ensures its robustness across diverse datasets and scenarios, offering a reliable evaluation measure Maini et al. (2024); Ma et al. (2024).
@@ 5.2.2 Logit-Based Evaluation
Membership Inference Attacks (MIAs) determine whether a specific data point is part of a model’s training data (member) or originates from outside the training set (non-member). Therefore, it serves as a tool to evaluate unlearning efficacy Jin et al. (2024); Yao et al. (2024).  Jin et al. [2024] used various MIA methods to assess the robustness of unlearning and found that many unlearning methods failed under such attacks.
5.2.3 Model Intervention
Relearning through Fine-tuning
Hu et al. [2024] suggested to use benign, publicly available data loosely related to the unlearned information. For example, while specific details about virus design might be unavailable, general knowledge such as “What is avian influenza?” is readily accessible. Their experiments demonstrate that augmenting with such benign data can inadvertently cause LLMs to relearn harmful knowledge, a finding further validated by Łucki et al. (2024).
LLM Quantization
facilitates faster inference and efficient deployment on resource-constrained devices with minimal performance degradation. However, recent studies Zhang et al. (2024c); Łucki et al. (2024) indicated that quantization can inadvertently aid in recovering unlearned knowledge. Zhang et al. [2024c] reported that for unlearning methods constrained by utility requirements, unlearned models retain an average of 21% of the intended forgotten knowledge in full precision, which dramatically increases to 83% after 4-bit quantization.
Activation Intervention
Recent research Arditi et al. (2024) demonstrated that refusal behavior in language models is mediated by a single direction in the activation space. Building upon this, outputs from each transformer block of both original and unlearned models are collected using the forget preference dataset Łucki et al. (2024); Seyitoğlu et al. (2024), and the refusal direction for each layer is computed by analyzing the differences in means. During inference, the refusal direction is removed at each layer. Extensive experiments confirmed the effectiveness of this approach.
@@ 6 Future Directions
Theoretical framework
@@
Unlearning in Multimodal Models
While numerous multimodal datasets have been introduced for multimodal unlearning, current methods remain largely confined to text-based unlearning approaches Ma et al. (2024); Liu et al. (2024b); Dontsov et al. (2024). Future research should prioritize the development of techniques capable of identifying and isolating modality-specific representations within MLLMs. Additionally, robust evaluation benchmarks are essential for assessing the effectiveness of multimodal unlearning methods in disentangling representations where knowledge is intertwined across both textual and visual inputs.
Real-World Complex Requests
Current unlearning methods primarily focus on removing specific data points from the model, requiring explicit target data points (sequences) to be provided. However, real-world unlearning requests may differ from this assumption. A significant future direction for LLM unlearning lies in addressing more complex requests, such as entity-level unlearning, which aims to remove all knowledge related to a specific entity across diverse contexts and associations. This involves not only forgetting explicit facts, but also erasing implicit or derived knowledge. Choi et al. [2024] introduced datasets to evaluate the effectiveness of algorithms in entity-level unlearning tasks. Looking ahead, even more complex scenarios may emerge, such as removing all information about a specific organization, or erasing entire domains of knowledge such as medical or criminal records.
@@ 7 Conclusion
We provided a comprehensive survey of recent advances in large language model (LLM) unlearning. We began by defining the problem and outlining the foundational settings of LLM unlearning. To offer a structured understanding, we proposed a novel taxonomy that categorizes existing research from diverse perspectives. We futher explored the methodologies used to implement unlearning and evaluates the effectiveness of these approaches in achieving the desired forgetting. Finally, we examined the key challenges in the field and identified promising directions for future research, thus offering valuable insights for researchers and practitioners.

### v2

@@ 3.3 Leveraging Auxiliary Models
$\min_{\theta_{a}}\mathcal{L}(\theta_{a})=\min_{\theta_{a}}\{\mathcal{L}_{f}(\theta_{a})-\beta\mathcal{L}_{r}(\theta_{a})\}.$ (10)
@@
$l(y|x;\theta_{ori})-\alpha\cdot{ReLU}(l(y|x;\theta_{a})-l(y|x;\theta_{ori}))$ (11)
@@
Task vectors
, defined as $\tau=\theta_{a}-\theta_{ori}$, steer the model behavior by editing the weight space, thus enabling operations such as negation and addition for applications such as unlearning and multi-task learning. Negating task vectors effectively suppress behaviors such as mitigating toxic language generation Ilharco et al. (2023). Liu et al. [2024d] enhanced fine-tuning with modules targeting harmful knowledge: guided distortion, random disassociation, and preservation divergence. Ethos Gao et al. (2024) distinguishes general and undesired knowledge by projecting task vectors onto principal components. By negating only undesired components, Ethos minimizes collateral damage to model utility, thus achieving unlearning.
@@ 3.4 Input/Output-based Unlearning
Method Utility Preservation Robust Forgetting Efficiency
Relabeling-Based Fine-Tuning Low Medium High
Gradient Ascent Low High High
Preference Optimization High Medium High
Reinforcement Learning High Medium High
Representation Engineering High Medium Medium
Locate-Then-Unlearn Methods High High Medium
Contrastive Decoding Medium Medium Low
Task Vector Low Medium Low
Knowledge Distillation Medium Medium Low
Input/Output-based Unlearning High Low High
Table 2: Comparison of various methods in terms of utility preservation, robust forgetting, efficiency.
Summary of Unlearning Methodologies
Table 2 provides a comparative overview of existing unlearning methodologies, assessing their strengths and limitations across three key dimensions: utility preservation, robust forgetting, and efficiency. These methods reflect varying design principles, leading to distinct trade-offs. For example, GA excels in robustly forgetting target data but significantly compromises utility, often resulting in degraded model performance. On the other hand, methods like PO and RL strike a more favorable balance, achieving high utility with moderate forgetting and efficiency, making them appealing for scenarios requiring minimal side effects. Locate-then-unlearn techniques stand out by offering both high utility preservation and strong forgetting capabilities, although they often involve computationally intensive attribution analysis. Approaches that rely on auxiliary models, such as contrastive decoding or knowledge distillation, tend to suffer from reduced efficiency due to added model complexity. Overall, the choice of method hinges on the specific unlearning goal—whether prioritizing forgetting effectiveness, maintaining model utility, or optimizing computational cost.
Benchmark Data Source # Volume Traditional Evaluation Adversarial Evaluation
Image Text
WHP Eldan et al. (2024) - Harry Potter books 330 GPT-Eval/Probability -
TOFU Maini et al. (2024) - Fictitious Profiles 4,000 Probability/Rouge/TR Paraphrasing
WMDP Li et al. (2024b) - Safety-related documents 3,668 Probe Accuracy Jailbreak Prompts
RWKU Jin et al. (2024) - Public figure knowledge 2,4510 Rouge Paraphrasing, Jailbreak Prompts,MIA
Łucki et al. (2024) - WMDP - Probe Accuracy Relearning,Intervention,Jailbreak Prompts,Pruning
Lynch et al. (2024) - WHP - GPT-Eval Relearning, Jailbreak Prompts
FIUBench Ma et al. (2024) Synthetic face images Generated private data 8,000 Rouge/GPT-Eval/TR/EM -
MMUBench Li et al. (2024a) Common visual concepts Related knowledge 2,000 Rouge/GPT-Eval/EM Paraphrasing, MIA, Jailbreak Prompts
MLLMU-Bench Liu et al. (2024b) Fictitious and celebrity faces Related profiles 20,754 Accuracy/Rouge Paraphrasing
CLEAR Dontsov et al. (2024) Synthetic face images Data from TOFU 7,700 TR, Probability, Rouge Paraphrasing
Table 3: Overview of existing benchmarks for LLM unlearning. Some Łucki et al. (2024); Lynch et al. (2024) focus on adversarial evaluation using existing datasets. TR: Truth Ration, EM: Exact Match.
@@ 4 Benchmarks
This section provides a detailed description of the commonly used benchmarks, addressing areas such as copyright, privacy, and AI alignment. Notably, WMDP and RWKU are designed for direct-unlearning, while other benchmarks are used within the fine-tuning-then-unlearning paradigm.
@@ 4.1 Unimodal Benchmarks
Who is Harry Potter (WHP) Eldan et al. (2024) evaluates unlearning of Harry Potter-related information using a dataset combining the original books (2.1M tokens) with synthetic content (1M tokens). Unlearning effectiveness is assessed through 330 Harry Potter-related questions scored with a GPT-4-based evaluation.
TOFU Maini et al. (2024) includes 200 synthetic author profiles with 20 question-answer examples each, ensuring no overlap with existing training data. The benchmark also includes 100 real-world author profiles and 117 world facts, comprehensively evaluating model utility after unlearning.
WMDP Li et al. (2024b) comprises 3,668 multiple-choice questions on biosecurity, cybersecurity, and chemical security, curated by experts to evaluate hazardous knowledge while excluding sensitive information. It serves as a benchmark for assessing LLMs’ hazardous knowledge and unlearning methods for AI alignment.
RWKU Jin et al. (2024) includes 200 unlearning targets centered on well-known public figures, comprising 13,131 multi-level forget probes and 11,379 neighbor probes. In addition, it incorporates a wide range of adversarial evaluation methods, including membership inference attacks (MIA), jailbreak prompts, and others.
@@ 5 Evaluation Measures
We group existing measures into two categories. Classical evaluation uses standard utility metrics to assess forgetting (via performance drop on the forget set) and utility preservation (on the retain set and world knowledge). Adversarial evaluation tests whether forgetting is robust or merely superficial suppression.
5.1 Classical Evaluation
To evaluate utility preservation and forgetting effectiveness, several classical metrics are commonly employed. (Probing) Accuracy assesses whether the unlearned model maintains its world knowledge without performance degradation. GPT-Eval uses large language models as evaluators to measure multiple dimensions beyond traditional metrics; for instance, Ma et al. (2024) employ GPT-4o Mini to score correctness, helpfulness, and relevance, producing an overall score between 0 and 1. ROUGE quantifies the similarity between generated outputs and ground truth responses Maini et al. (2024); Yuan et al. (2024). Probability-based evaluation estimates the model’s confidence by computing the normalized conditional likelihood of the target output. Truth Ratio is specially proposed to compare the model’s likelihoods of correct versus incorrect answers for a given question and has been used in many benchmarks Maini et al. (2024); Ma et al. (2024); Dontsov et al. (2024). Specifically, it is the ratio of the average normalized conditional probability of perturbed incorrect answers $\bar{y}\in\mathcal{A}$ to that of a paraphrased correct answer $\tilde{y}$. Both $\bar{y}$ and $\tilde{y}$ can be generated by LLMs. A lower truth ratio indicates better forgetting. It is defined as:
$R_{\text{truth}}(y)=\frac{\frac{1}{|\mathcal{A}|}\sum_{\bar{y}\in\mathcal{A}}p(\bar{y}|x;\theta_{\text{unl}})^{1/|\bar{y}|}}{p(\tilde{y}|x;\theta_{\text{unl}})^{1/|\tilde{y}|}}$
On the retain and world knowledge sets, $\max(0,1-R_{\text{truth}}(y))$ is used to measure how well the model preserves relevant information.
5.2 Adversarial Evaluation
Recent work Lynch et al. (2024); Ma et al. (2024) emphasize the essence of distinguishing real forgetting from mere suppression, where suppressed knowledge can be recovered via adversarial techniques. This motivates the summation of adversarial evaluation to capture such subtle distinctions.
Input-based evaluation
assesses whether a model has truly forgotten information by modifying the input rather than probing its internal representations. Paraphrasing involves rephrasing questions or translating them into other languages (e.g., Spanish or Russian) to test whether the model still recalls unlearned content Maini et al. (2024); Lynch et al. (2024). Jailbreak prompts attempt to revive forgotten knowledge by providing contextual cues or demonstrations during inference Lynch et al. (2024), or by crafting adversarial inputs. These include both black-box strategies (e.g., role-playing) and white-box methods (e.g., optimized prefixes) to bypass unlearning and extract suppressed information Łucki et al. (2024).
Logit-based evaluation
@@ 5.2 Adversarial Evaluation
assesses unlearning effectiveness by analyzing changes in the model’s output distributions—typically probabilities or logits—before and after unlearning. Kolmogorov-Smirnov Test (KS-Test) measures the divergence between output distributions by comparing their cumulative distribution functions (CDFs). A lower KS statistic indicates greater unlearning success. Its non-parametric nature makes it robust across various datasets and tasks Maini et al. (2024); Ma et al. (2024). Moreover, Membership Inference Attacks (MIAs) determine whether a specific data point is part of a model’s training data (member) or originates from outside the training set (non-member). Hence, it is applied to evaluate unlearning efficacy Jin et al. (2024); Yao et al. (2024).  Jin et al. [2024] used various MIA methods to assess the robustness of unlearning and found that many unlearning methods failed under such attacks.
Model intervention methods
evaluate the robustness of unlearning by directly modifying the model’s parameters, activations, or numerical precision—interventions that may inadvertently reverse unlearning and expose residual memorization. Relearning through fine-tuning refers to the phenomenon where continued training on benign and loosely related data (e.g., "What is avian influenza?") causes the model to recover previously forgotten knowledge. This suggests that unlearning may suppress rather than fully remove the underlying representations Hu et al. (2024); Łucki et al. (2024). LLM quantization offers another perspective, where reducing the model’s precision—such as converting weights to 4-bit—can increase the likelihood of forgotten content re-emerging, thereby weakening the unlearning effect Zhang et al. (2025); Łucki et al. (2024). Similarly, Łucki et al. (2024) evaluate neuron pruning as a means of assessing residual memorization. Finally, activation intervention techniques analyze the model’s internal activations to identify and remove the so-called refusal direction—a vector derived from differences between the original and unlearned models. Suppressing this direction reduces refusal behavior and enables the model to regenerate responses that were assumed to be forgotten Arditi et al. (2024); Łucki et al. (2024); Seyitoğlu et al. (2024); Li et al. (2025b).
@@ 6 Future Directions
Theoretical frameworks are methodologically demanding.
@@
Multimodal unlearning shows promising potential.
While numerous multimodal datasets have been introduced for multimodal unlearning, current methods remain largely confined to text-based unlearning approaches Ma et al. (2024); Liu et al. (2024b); Dontsov et al. (2024). Future research should prioritize the development of techniques capable of identifying and isolating modality-specific representations within MLLMs. Moreover, robust evaluation benchmarks are essential for assessing the effectiveness of multimodal unlearning methods in disentangling representations where knowledge is intertwined across both texts and images.
Real-world complexity is crucial for robust evaluation.
Current unlearning methods primarily focus on removing specific data points from the model, requiring explicit target data points (sequences) to be provided. However, real-world unlearning requests may differ from this assumption. A significant future direction for LLM unlearning lies in addressing more complex requests, such as entity-level unlearning, which aims to remove all knowledge related to a specific entity across diverse contexts and associations. This involves not only forgetting explicit facts but also erasing implicit or derived knowledge. Choi et al. [2024] introduced datasets to evaluate the effectiveness of algorithms in entity-level unlearning tasks. Looking ahead, even more complex scenarios may emerge, such as removing all information about a specific organization, or erasing entire domains of knowledge, such as medical or criminal records.
@@ 7 Conclusion
We provided a comprehensive survey of recent advances in LLM unlearning. We began by defining the problem and outlining the foundational settings of LLM unlearning. To offer a structured understanding, we proposed a novel taxonomy that categorizes existing research from diverse perspectives. We further explored the methodologies used to implement unlearning and evaluates the effectiveness of these approaches in achieving the desired forgetting. Finally, we examined the key challenges in the field and identified promising directions for future research, thus offering valuable insights for researchers and practitioners.
Limitations
This survey mainly has the following limitations:
No experimental benchmarks
Without original experiments, this paper cannot offer empirical validation of the theories or concepts. This limits the paper’s ability to contribute new, verified knowledge to the field.
Potential omissions
We have made our best effort to compile the latest advancements. Due to the rapid development in this field, there is still a possibility that some important work may have been overlooked.
Ethics and Broader Impact
We anticipate no significant ethical concerns in our work. As a survey of recent progress in this research area, our study does not involve experimental implementation, the use of sensitive datasets, or the employment of annotators for manual labeling.